(2) Consider the sample: 9, 6. 7, 9, 10, 8 a) Find: the mean, the median, the mode b) Find: the range, the quartiles Q1, Q3, the variance, the standard deviation c) Draw a box-plot for the given data set.
Answer #1

a)

sorted data

6,7,8,9,9,10

mean = 8.16666

median =   (8+9)/2 = 8.5

mode = 9 as it occurs most of time

b)

range= maximum - minimum = 10-6 = 4

Q1 = median for first half

= 7

Q3 = median of last half = 9

Step 1: Find the mean. X In this example: X- 8.1667. (use this calculator for step by step explanation on how to find mean). Step 2: Create the following table data data-mean 2.1667 1.1667 0.1667 0.8333 0.8333 1.8333 (data - mean) 4.69458889 1.36118889 0.02778889 0.69438889 0.69438889 3.36098889 6 10 Step 3: Find the sum of numbers in the last column to get. Ci - X) 10.8333 Step 4: Calculate variance using the above formula *i-X10.8333 uar- 2.1667 6-1

variance = 2.16666

sd = SQRT(VARIANCE) = 1.47196

c)

Box Plot Chart 16.5 14.5 12.5 10.5 8.5 6.5 2.5 Sanple 1

Box Plot Chart 16.5 14.5 12.5 10.5 8.5 6.5 2.5 Sanple 1